how to calculate pv of cash flow
cumulative pv factor for 12 periods
and pv factor at nth period
Answer #1

1: PV of cash flows = Sum of discounted value of cash flows

=  CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

2: CumulativePV factor for 12 periods = (1-1/(1+rate)^number of terms)/rate

=(1-1/(1+rate)^12)/rate

3: PV factor at nth period = 1/(1+Rate)^n